# 4Université de Strasbourg, CNRS UMR 7550, Observatoire astronomique de Strasbourg, 67000 Strasbourg, France 5DIO, Observatoire de Paris, CNRS, PSL, 75014, Paris, France 6IDRIS, CNRS, F-91403 Orsay, France

### v1

@@ 1 Introduction
The present paper is the second in a series introducing YOLO-CIANNA, a deep learning regression-based method for source detection and characterization. The series aims to evaluate the method’s performance on simulated data from the SDCs, in preparation for its applications to observational data from SKA precursors and pathfinders. The first paper, Cornu et al. (2024) hereafter Paper I, covers the full method description and its application to simulated continuum images from the SDC1 dataset (Bonaldi et al., 2021). It also includes a concise review of ML object detection methods along with some examples of astronomical applications. The method description emphasizes specific design choices tailored to astronomical datasets. This first paper concludes with a discussion of the biases and limitations associated with the SDC1 challenge definition, and presents possible improvements to the method itself.
In the present work, we extend the results of Paper I by generalizing YOLO-CIANNA to 3D source detection in hyperspectral data and apply it to simulated HI emission cubes from the SDC2. The method was initially developed and used by team MINERVA (Machine Learning for Radioastronomy at Observatoire de Paris) for its participation in the SDC2, in which the team secured first place. Therefore, this paper provides a detailed description of an updated version of MINERVA’s method.
@@ 2.1 Data cubes and source catalogs
The SDC2 is a galaxy detection and characterization challenge in simulated 3D HI cubes representative of future SKA-mid observations. A complete description of the challenge and of the associated data can be found in the SDC2 summary paper Hartley et al. (2023), which details the simulation of the source catalog using TRECS (Bonaldi et al., 2019), the construction of the sky model, and the addition of instrumental effects. The primary data product is a MAIN cube covering a 20 square degree region with a synthesized beam of 7 arcsec and a pixel resolution of 2.8 arcsec. It covers a frequency bandwidth of 950-1150 MHz (z=0.235z=0.235-0.495) sampled at 30 kHz. This results in a massive cube (851 GB) of size $5851\times 5851\times 6668$ [pixels, pixels, channels], corresponding to $228\times 10^{9}$ voxels, a voxel being defined as the volume of a square pixel over one frequency channel. Continuum images of the same field are also provided with an identical spatial resolution and bandwidth, but sampled at a 10 MHz resolution. Thermal noise consistent with 2000h of SKA integration time was added, yielding 26 to 31 µJy $\mbox{beam}^{-1}$ per channel. The cube also features continuum subtraction residues, instrumental artifacts, and radio frequency interferences (RFI) subtraction residues, improving its realism.
@@
Figure 1 shows a cutout around one of the brightest sources from the truth catalog. To improve contrast, voxel values are rescaled following $p^{\prime}_{i}=\tanh\left(3p_{i}/\max_{p}\right)$, where $p_{i}$ and $p^{\prime}_{i}$ are the original and rescaled voxel values, respectively, and $max_{p}$ is the maximum voxel value in the cutout. To better visualize the inner 3D structure, we set the voxel transparency as inversely proportional to their rescaled value. This figure illustrates the typical 3D structure of a galaxy observed in HI, where the line width correlates with its rotation. With the SDC2 observational setup, sources are compact in the plane of the sky but extended over tens of channels, with less signal at the central coordinates. The signal is mostly aligned in a plane, which defines the source angles, $PA$ and $i$.
@@ 2.2 SDC2 scoring metric
In addition, we can also estimate mean subscores and a global average characterisation score as
@@ 2.3 Detectability selection function
YOLO-CIANNA is a supervised ML approach and therefore requires a set of labeled examples. For this, we rely on the provided LDEV cube and truth source catalog. However, as with the SDC1, the vast majority of the listed sources are too faint to be detectable ($\sim 85\%$). Including undetectable sources in the target catalog may be debatable from a challenge perspective. Still, the presence of faint sources in the simulation enhances background noise realism. It also ensures a realistic source signal-to-noise ratio (S/N) distribution in the target catalog without imposing an arbitrary detectability cut.
As with most supervised ML approaches, our method is sensitive to mislabeling in the training examples. It involves objects that should be detectable but not labeled as targets, or labeled targets that are too faint to be distinguishable from the background (Sects. 2.2 and 3.4 of Paper I). Therefore, we define a detectability selection function to curate the LDEV truth catalog. This function combines the integrated line flux $f$ with an estimate of the source volume brightness $V_{b}$. We approximate the source volume by a cylinder with a radius equal to its beam-convolved HI size, expressed in pixels, and a length equal to its line width, in channels. The volume brightness is obtained as
@@
where $\hat{S}_{deg}$ is the target source size in degree, $B_{s}$ the synthesized beam size in degree, $P_{s}$ the pixel size in degree, $\hat{w}_{20}$ the line width in km s-1, $c$ the speed of light in km s-1, $\hat{\nu}$ the target central frequency in Hz, $\nu_{HI_{0}}$ the rest frequency of HI in Hz, and $C_{s}$ the frequency channel size in Hz. The resulting volume brightness is in Jy Hz voxel-1. The selection function is then defined as
@@
Figure 2: Top: Two-dimensional histogram of LDEV truth source volume as a function of their integrated line flux. The red dashed line represents our selection function, with sources to the right of the line being selected. Bottom: line flux distribution for the full LDEV catalog and the selected sample.
@@ 2.4 Data cube preprocessing
We perform a series of preprocessing steps to enhance source detectability and prepare the data cube for use with our detector. We start by identifying the sky positions for which the mean signal over all continuum images is above $3\times 10^{-3}$ $\mbox{Jy}\,\mbox{beam}^{-1}$ and set all the corresponding cube voxels to zero, masking continuum-subtraction residuals. We then normalize every frequency channel by its standard deviation. This reduces the impact of RFI subtraction residuals and makes the channel noise homogeneous across the whole bandwidth. It also allows us to consider the detection task as invariant with the frequency position. This way, a detector can be trained and applied over subcubes along the frequency axis.
Due to the high-frequency resolution, the source signal is diluted over multiple channels. This could be mitigated by smoothing the cube along the frequency axis or averaging every few channels, which would increase apparent S/N for some sources. However, it would destroy part of the fine-grained 3D structure that the detector could exploit. We observed that such preprocessing allows for faster training convergence but reduces the final score. We therefore retain the original frequency resolution to maximise detection performance.
The last step is to renormalize the cube. As for the SDC1, we apply a scaled hyperbolic tangent transformation to the voxels following $p_{i}^{\prime}=\tanh{\left(\alpha p_{i}\right)}$ where $p_{i}$ is the cube voxel value after the previous preprocessing steps, and $p_{i}^{\prime}$ is the final normalized voxel value. The parameter $\alpha$ controls the steepness of the curve and thus impacts the final signal contrast and saturation. After some exploration, we adopted $\alpha=0.3$ for the SDC2. We normalize both the MAIN and LDEV cubes once and store them in a reduced unsigned 16-bit integer format for faster dynamic loading. We observed no impact of this reduction on the final score, provided the reduction is done after the tanh normalization. To further reduce the loading time, we split the MAIN cube into chunks of $512\times 512\times 6668$ (3.5 GB), cutting along the two sky axes. The cuts are made based on our prediction pipeline setup (Sect. 3.5), ensuring partial overlap and padding when necessary. The LDEV cube is left as a single file and fully loaded into system memory for the training phase (Sect. 3.3.1).
@@ 3.1 YOLO-CIANNA generalization for 3D detection
In Paper I (Cornu et al., 2024), we provide a complete description of the YOLO-CIANNA method for detecting, classifying, and characterizing objects in 2D images. Here, we present adaptations of the formalism for detecting 3D objects. As explained in Paper I, YOLO-CIANNA is a regression-based detection method implemented as part of the broader CIANNA development framework (Cornu, 2024) [^2] CIANNA is open-source and freely accessible through GitHub https://github.com/Deyht/CIANNA. The version used in this paper corresponds to the V-1.0 release 10.5281/zenodo.12806325.. The YOLO-CIANNA method constrains a fully convolutional neural network (CNN) backbone to create a mapping from an input data space to a regular grid of detection units. Each unit is associated with an independent subregion of the input space and is tasked with detecting all objects for which the central coordinate lies within this region. Objects are represented as bounding boxes. During the training phase, the network is fed with labeled images. Each detection unit produces a set of candidate boxes that are matched and compared with the known list of target boxes. This comparison is then used to adjust network weights so future predictions will be closer to the target. Additionally, the model learns to predict an objectness score for each detection, representing its confidence that a given predicted box corresponds to a real object.
To generalize the formalism presented in Paper I toward 3D detection, we add two new dimensions to the detection units’ output vector to predict both the central position $z$ and the size $d$ of a box along the third dimension. We express a 3D bounding box as a function of output vector elements, following
@@
We also update the box proximity and similarity metric used in many parts of the method. For this, we rely on the Intersection over Union (IoU) measurement generalized to 3D boxes. In this paper, we use its distance-aware variant, the $\rm DIoU$} (Zheng et al., 2020), bounded within [1,1][-1,1], and defined as
@@ 3.2 Network backbone architecture
Using the updated method, a classical 2D CNN backbone could be used for 3D detection by treating the third dimension as a stack of independent 2D input channels. However, taking advantage of spatial invariance through fine-grained gridding across all positional dimensions usually improves detection accuracy. Thus, we grid the third dimension into independent detection units as well. For this, we rely on CIANNA’s ability to build 3D spatial convolutional layers and design a fully 3D network backbone. The output layer is therefore directly in the form of a 3D output grid (indexed by $g^{x}$, $g^{y}$, $g^{z}$), and detection units are structurally centered on their associated input subvolume. This structure also reduces the number of learnable weights required in the model compared to a 2D backbone approach. More details about the interplay between our method and the network backbone can be found in Sect. 3.6 and Appendix A of Paper I.
@@
Even though a fully convolutional architecture is not bound to a specific input size, we chose to use a fixed subcube input dimension of $64\times 64\times 256$. The ratio between the spatial and spectral dimensions has been defined regarding the typical source shape (Fig. 1) and constraints in both the training and prediction pipelines (Sects. 3.3.1 and  3.5). In the SDC2 data, and likely for any hyperspectral data as well, the signal distribution along the frequency dimension is very different from that of the two sky dimensions. Although all three axes are considered spatial in terms of the box definition and matching criteria, we use anisotropic filters and stride configurations to account for the intrinsic differences in signal distribution across dimensions. Our best-performing SDC2 backbone is presented in Fig. 3, which was obtained through an extensive search over the architecture and hyperparameter space using our SDC1 backbone as a starting point. The detailed configuration of our backbone architecture can be found in the example scripts hosted on CIANNA’s GitHub repository and archived at xx.xxxx/zenodo.xxxxxxxx [^3] Pre-review archive at https://share.obspm.fr/s/swyCT7BgEGjtZK3. As we discussed in Appendix B of Paper I, we observed that classical backbones constrained from generic image datasets underperformed on the SDC2 task, and even more if used pre-trained.
@@
Despite being deeper with 23 convolutional layers, this architecture only contains $3.35$ million learnable weights, which is a quarter of what we had in our 17-layer SDC1 backbone. This results from a reduction in the number of filters, which is motivated by the reduced number of sources in the training catalog (1383 sources, Sect. 2.3). Still, the lesser morphological diversity in simulated HI sources compared to simulated continuum sources allows the network to achieve high detection and characterization accuracy with fewer parameters. We note that the first layer convolves the frequency axis with filters spanning height channels with a stride of two, which searches for large spectral structures and reduces the frequency axis dimension, compensating for the absence of frequency smoothing in the data preparation. Like our SDC1 backbone, the first few layers progressively reduce both the spatial dimensionality and the number of filters, acting primarily as a small denoising subnetwork tailored to this detection task. We also include the typical darknet sequence of alternating $3\times 3$ convolutions with $1\times 1$ convolutions over a smaller number of filters (Redmon & Farhadi, 2017). This forces the network to find more compressed representations between spatial pattern extractions steps, which reduces the number of learnable weights compared to stacking identically configured spatial convolutions. Here, these compression steps have a secondary purpose, which is to capture spectral patterns with 3D filters of size $1\times 1\times 3$. This increases the receptive field over the spectral dimension faster than for the two sky dimensions, which helps extract extended patterns along the frequency axis. All layers except the last one use a leaky-ReLU activation with a 0.1 leakage factor. Like our SDC1 backbone, we have a group normalization layer between C20 and C21 with a group size of two (Wu & He, 2018), and a 25% dropout rate on the C21 layer (Srivastava et al., 2014), both for regularization purposes.
@@
Even though we could reduce the spectral dimension to have a regular output grid in all dimensions, we achieve better spectral localisation by sampling it over a finer output detection grid. The total reduction factors for each dimension at the end of the network are then $8:8:16$, and each detection unit is responsible for an $8\times 8\times 16$ input subvolume. The receptive field at the last layer is $66\times 66\times 294$, which corresponds to the input subvolume accessible to each detection unit to make a decision. Removing layers significantly reduces the achievable score, indicating that network depth compensates for fewer learnable weights in achieving the required expressivity for the task.

### v2

@@ 1 Introduction
The present paper is the second in a series introducing YOLO-CIANNA, a deep-learning regression-based method for source detection and characterization. The series aims to evaluate the method’s performance on simulated data from the SDCs, in preparation for its applications to observational data from SKA precursors and pathfinders. The first paper, Cornu et al. (2024) hereafter Paper I, covers the full method description and its application to simulated continuum images from the SDC1 dataset (Bonaldi et al., 2021). It also includes a concise review of ML object detection methods along with some examples of astronomical applications. The method description emphasizes specific design choices tailored to astronomical datasets. This first paper concludes with a discussion of the biases and limitations associated with the SDC1 challenge definition, and presents possible improvements to the method itself.
In the present work, we extend the results of Paper I by generalizing YOLO-CIANNA to 3D source detection in hyperspectral data and apply it to simulated HI emission cubes from the SDC2. The method was initially developed and used by team MINERVA (Machine Learning for Radioastronomy at Observatoire de Paris) for its participation in SDC2, in which the team secured first place. Therefore, this paper provides a detailed description of an updated version of MINERVA’s method.
@@ 2.1 Data cubes and source catalogs
The SDC2 is a galaxy detection and characterization challenge in simulated 3D HI cubes representative of future SKA-mid observations. A complete description of the challenge and of the associated data can be found in the SDC2 summary paper Hartley et al. (2023), which details the simulation of the source catalog using TRECS (Bonaldi et al., 2019), the construction of the sky model, and the addition of instrumental effects. The primary data product is a MAIN cube covering a 20 square degree region with a synthesized beam of 7 arcsec and a pixel resolution of 2.8 arcsec. It covers a frequency bandwidth of 950-1150 MHz (z=0.235z=0.235-0.495) sampled at 30 kHz. This results in a massive cube (851 GB) that is $5851\times 5851\times 6668$ (pixels, pixels, channels) in size, corresponding to $228\times 10^{9}$ voxels, a voxel being defined as the volume of a square pixel over one frequency channel. Continuum images of the same field are also provided with identical spatial resolution and bandwidth, but sampled at a 10 MHz resolution. Thermal noise consistent with 2000h of SKA integration time was added, yielding 26 to 31 µJy $\mbox{beam}^{-1}$ per channel. The cube also features continuum subtraction residues, instrumental artifacts, and radio frequency interference (RFI) subtraction residues, improving its realism.
@@
Figure 1 shows a cutout around one of the brightest sources from the truth catalog. To improve contrast, voxel values have been rescaled following $p^{\prime}_{i}=\tanh\left(3p_{i}/\max_{p}\right)$, where $p_{i}$ and $p^{\prime}_{i}$ are the original and rescaled voxel values, respectively, and $max_{p}$ is the maximum voxel value in the cutout. To better visualize the inner 3D structure, we set voxel transparency to be inversely proportional to their rescaled value. This figure illustrates the typical 3D structure of a galaxy observed in HI, where the line width correlates with its rotation. With the SDC2 observational setup, sources are compact in the plane of the sky but extended over tens of channels, with lower signal intensity at the central coordinates. The signal is mostly aligned in a plane, which defines the source angles, $PA$ and $i$.
@@ 2.2 SDC2 scoring metric
In addition, we can also estimate mean subscores and a global average characterization score as
@@ 2.3 Detectability selection function
YOLO-CIANNA is a supervised ML approach and therefore requires a set of labeled examples. For this, we relied on the provided LDEV cube and truth source catalog. However, as with the SDC1, the vast majority of the listed sources are too faint to be detectable ($\sim 85\%$). Including undetectable sources in the target catalog may be debatable from a challenge perspective. Still, the presence of faint sources in the simulation enhances background noise realism. It also ensures a realistic source signal-to-noise ratio (S/N) distribution in the target catalog without imposing an arbitrary detectability cut.
As with most supervised ML approaches, our method is sensitive to mislabeling in the training examples. It involves objects that should be detectable but not labeled as targets, or labeled targets that are too faint to be distinguishable from the background (Sects. 2.2 and 3.4 of Paper I). Therefore, we defined a detectability selection function to curate the LDEV truth catalog. This function combines the integrated line flux, $f$, with an estimate of the source volume brightness, $V_{b}$. We approximate the source volume by a cylinder with a radius equal to its beam-convolved HI size, expressed in pixels, and a length equal to its line width, in channels. The volume brightness was obtained as
@@
where $\hat{S}_{deg}$ is the target source size in degree, $B_{s}$ the synthesized beam size in degree, $P_{s}$ the pixel size in degree, $\hat{w}_{20}$ the line width and $c$ the speed of light in kilometer per second, $\hat{\nu}$ the target central frequency in Hertz, $\nu_{HI_{0}}$ the rest frequency of HI in Hertz, and $C_{s}$ the frequency channel size in Hertz. The resulting volume brightness is in Jansky Hertz voxel-1. The selection function is then defined as
@@
Figure 2: Top: Two-dimensional histogram of LDEV truth source volume as a function of their integrated line flux. The dashed red line represents our selection function, with sources to the right of the line being selected. Bottom: Line flux distribution for the full LDEV catalog and the selected sample.
@@ 2.4 Data cube preprocessing
We performed a series of preprocessing steps to enhance source detectability and prepared the data cube for use with our detector. We started by identifying the sky positions for which the mean signal over all continuum images is above $3\times 10^{-3}$ $\mbox{Jy}\,\mbox{beam}^{-1}$ and set all the corresponding cube voxels to zero, masking continuum-subtraction residuals. We then normalized every frequency channel by its standard deviation. This reduces the impact of RFI subtraction residuals and makes the channel noise homogeneous across the whole bandwidth. It also allows us to consider the detection task as invariant with respect to the frequency position. This way, a detector can be trained and applied over subcubes along the frequency axis.
Due to the high frequency resolution, the source signal is diluted over multiple channels. This could be mitigated by smoothing the cube along the frequency axis or averaging every few channels, which would increase the apparent S/N for some sources. However, it would destroy part of the fine-grained 3D structure that the detector could exploit. We observed that such preprocessing allows for faster training convergence but reduces the final score. We therefore retained the original frequency resolution to maximize detection performance.
The last step was to renormalize the cube. As for the SDC1, we applied a scaled hyperbolic tangent transformation to the voxels following $p_{i}^{\prime}=\tanh{\left(\alpha p_{i}\right)}$, where $p_{i}$ is the voxel value after the previous preprocessing steps, and $p_{i}^{\prime}$ is the final normalized voxel value. The parameter $\alpha$ controls the steepness of the curve and thus impacts the final signal contrast and saturation. After some exploration, we adopted $\alpha=0.3$ for the SDC2. We normalized both the MAIN and LDEV cubes once and stored them in a reduced unsigned 16-bit integer format for faster dynamic loading. We observed no impact of this reduction on the final score, provided the reduction is done after the tanh normalization. To further reduce the loading time, we split the MAIN cube into chunks of $512\times 512\times 6668$ (3.5 GB), cutting along the two sky axes. The cuts were made based on our prediction pipeline setup (Sect. 3.6), ensuring partial overlap and padding when necessary. The LDEV cube was left as a single file and fully loaded into system memory for training (Sect. 3.4).
@@ 3.1 YOLO-CIANNA generalization for 3D detection
In Paper I (Cornu et al., 2024), we provide a complete description of the YOLO-CIANNA method of detecting, classifying, and characterizing objects in 2D images. Here, we present adaptations of the formalism for detecting 3D objects. As is explained in Paper I, YOLO-CIANNA is a regression-based detection method implemented as part of the broader CIANNA development framework (Cornu, 2024) [^2] CIANNA is open-source and freely accessible through GitHub https://github.com/Deyht/CIANNA. The version used in this paper corresponds to the V-1.0 release 10.5281/zenodo.12806325.. The YOLO-CIANNA method constrains a fully convolutional neural network (CNN) backbone to create a mapping from an input data space to a regular grid of detection units. Each unit is associated with an independent subregion of the input space and is tasked with detecting all objects whose central coordinate lies within that region. Objects are represented as bounding boxes. During the training phase, the network is fed with labeled images. Each detection unit produces a set of candidate boxes, which are compared to the known list of target boxes through a matching function. The network weights are then adjusted to minimize the observed difference between its prediction and the expected result. Additionally, the model learns to predict an objectness score for each detection, representing its confidence that a given predicted box corresponds to a real object.
To generalize the formalism presented in Paper I to 3D detection, we added two new dimensions to the detection units’ output vector to predict both the box’s central position, $z$, and its size, $d$, along the third dimension. We express a 3D bounding box as a function of output vector elements, following
@@
We also updated the box proximity and similarity metric used in many parts of the method. For this, we relied on the intersection over union (IoU) measurement generalized to 3D boxes. In this paper, we use its distance-aware variant, the $\rm DIoU$} (Zheng et al., 2020), bounded within [1,1][-1,1], and defined as
@@ 3.2 Network backbone architecture
Using the updated method, a classical 2D CNN backbone could be used for 3D detection by treating the third dimension as a stack of independent 2D input channels. However, taking advantage of spatial invariance through fine-grained gridding across all positional dimensions usually improves detection accuracy. Thus, we gridded the third dimension into independent detection units as well. For this, we relied on CIANNA’s ability to build 3D spatial convolutional layers and design a fully 3D network backbone. The output layer is therefore directly in the form of a 3D output grid (indexed by $g_{x}$, $g_{y}$, $g_{z}$), and detection units are structurally centered on their associated input subvolume. This structure also reduces the number of learnable weights required in the model compared to a 2D backbone approach. More details about the interplay between our method and the network backbone can be found in Sect. 3.6 and Appendix A of Paper I.
@@
Even though a fully convolutional architecture is not bound to a specific input size, we chose to use a fixed subcube input dimension of $64\times 64\times 256$. The ratio between the spatial and spectral dimensions has been defined regarding the typical source shape (Fig. 1) and constraints in both the training and prediction pipelines (Sects. 3.4 and 3.6). In the SDC2 data, and likely for any hyperspectral data as well, the signal distribution along the frequency dimension is very different from that of the two sky dimensions. Although all three axes are considered spatial for the box definition and matching criteria, we used anisotropic filters and stride configurations to account for the intrinsic differences in signal distribution across dimensions. Our best-performing SDC2 backbone, presented in Fig. 3, was obtained through an extensive search over the architecture and hyperparameter space using our SDC1 backbone as a starting point. The detailed configuration of our backbone architecture can be found in the training scripts archived at 10.5281/zenodo.18403011 (Sect. 7). As we observed for SDC1 (Appendix B of Paper I), classical backbones developed for generic image datasets underperform when trained on the SDC2 task in comparison with a custom architecture. Using a pre-trained generic model as a starting point for training on the SDC2 task further degrades performance.
@@
Despite being deeper with 23 convolutional layers, this new 3D architecture contains only $3.35$ million learnable weights, which is a quarter of what we had in our 17-layer SDC1 backbone. This mainly results from a reduction in the number of filters per layer, motivated by the reduced number of sources in the training catalog (1383, Sect. 2.3). Still, the lesser morphological diversity in simulated HI sources compared to simulated continuum sources enables the network to achieve high detection and characterization accuracy with fewer learnable parameters. We note that the first layer convolves the frequency axis with filters spanning height channels, with a stride of two. This operation searches for large spectral structures and reduces the frequency axis dimension, compensating for the absence of frequency smoothing in the data preparation. Like our SDC1 backbone, the first few layers progressively reduce both the spatial dimensionality and the number of filters, acting primarily as a small denoising subnetwork tailored to this detection task. We also include the typical darknet sequence of alternating $3\times 3$ convolutions with $1\times 1$ convolutions over a smaller number of filters (Redmon and Farhadi, 2017). This forces the network to learn more compressed representations between spatial pattern extraction steps, reducing the number of learnable weights compared to stacking identically configured spatial convolutions. Here, these compression steps have a secondary purpose, which is to capture spectral patterns with 3D filters of size $1\times 1\times 3$. This increases the receptive field along the spectral dimension faster than for the two sky dimensions, allowing the identification of extended patterns along the frequency axis. All layers except the last one use a leaky-ReLU activation with a 0.1 leakage factor. As with our SDC1 backbone, we have a group normalization layer between C20 and C21 with a group size of two (Wu and He, 2018) and a 25% dropout rate on the C21 layer (Srivastava et al., 2014), which helps regularize the model.
@@
Even though we could reduce the spectral dimension to obtain a 3D output grid of equal size across all dimensions, we observed that sampling the spectral domain over more grid elements grants better spectral localization. The total reduction factors for each dimension at the end of the network are then $8:8:16$, and each detection unit is responsible for an $8\times 8\times 16$ input subvolume. The receptive field at the last layer is $66\times 66\times 294$, which corresponds to the input subvolume accessible to each detection unit for making a prediction. Removing layers anywhere in the final backbone measurably reduces the achievable score, while increasing the number of filters per layer leads to early overtraining. This likely indicates that the number of parameters is limited by the training sample size, and that network depth helps reach the required expressivity with fewer learnable parameters.